OpenStudy (anonymous):

Oh my! This next one is a rough!! the yearly production of a 6 foot orange tree is 20 lbs. of oranges. A 16ft. tree produces 94 lbs. let h represent the height of an orange tree and p the number of lbs. of oranges produced. Use the information to answer the question. list the ordered pairs what is the ordered pair for the first tree?

OpenStudy (amistre64):

\[\begin{array}c &h&p\\&--&--\\1:&6&20\\2:&16&93\end{array}\]

OpenStudy (amistre64):

such beautiful penmenship and I mess up the 94.... ack!!

OpenStudy (anonymous):

Wow! I don't understand that at all! Is that the answer?

OpenStudy (anonymous):

Are you there?

OpenStudy (amistre64):

the answer is (6,20)

OpenStudy (anonymous):

Thank you. thank you, thank you!!!

OpenStudy (amistre64):

the next ordered pair is (16,94) :)

OpenStudy (amistre64):

and if my psychic powers are up to date; the line they want is 16 94 -6 -20 ------ 10 74; slope = 74/10 y = (74/10)(x+6)+20 or some such ;)

OpenStudy (anonymous):

the next part of the question is find the point-slope form of an equation of the line through the two ordered pairs.

OpenStudy (amistre64):

point slope .... i was close :) y-20 = 74/10(x-6)

OpenStudy (anonymous):

p-20=37/5_h-6) or p-20=-37/5(h-6) or p=37/5h-122/5 or p-6=5/37(h-20)

OpenStudy (anonymous):

Can you help me with this part?